OpenStudy (anonymous):

how do you solve (x+1)^-1=x^-1+x graphically by converting it to an equivalent equation and how do you find the x-intercept

OpenStudy (zehanz):

Here's a graph drawn in Geogebra. Zooming in will do the trick! (Or do you have to do it with a calculator like the TI-84?) BTW there is no x-intercept...

OpenStudy (zehanz):

My bad, here's the graph:

OpenStudy (anonymous):

how would you convert the equation to an equivalent one

OpenStudy (mertsj):

\[\frac{1}{x+1}=\frac{1}{x}+x\]

OpenStudy (mertsj):

Multiply each term by the common denominator which is x(x+1)

OpenStudy (zehanz):

The question looks a bit cryptic to me: if you have to solve it graphically, why bother converting it to an equivalent equation?

OpenStudy (anonymous):

to question says.... solve the equation graphically by converting it to an equivalent equation with 0 on the right hand side and then find the xintercepts

OpenStudy (zehanz):

OK, if you do what @Mertsj said, you'll get a 3rd degree equation, that can only be solved graphically. Makes sense ;)